OpenStudy (anonymous):

Find the total length of the cardioid: r = 4 (1 + cos x) Help please

OpenStudy (anonymous):

Cardiods in polar form look like: \[r = 2a(1 + \cos\theta)\] The length is 16a

OpenStudy (anonymous):

need the working :/

OpenStudy (anonymous):

depending on what math you're in, the arc length of any curve defined in polar coordinates is given by \[\int\limits_a^b \sqrt{r^2+\left(\frac{dr}{d\theta}\right)^2} d\theta\]